Name: Daniel Finn

Content and Structure – The paper

1. Has a clearly stated topic or question that it will address – 10

2. Does address the topic or question – 5

3. Has a clearly articulated and logical structure:

a. First paragraph lays out the topic or question – 10 b. Later paragraphs have topic sentences and clearly identifiable main ideas – 10 c. Transitions between paragraphs are logical – 10

4. Accurately presents facts and viewpoints of sources – 10

5. Provides sufficient textual evidence to support the argument – 10

6. Is concise, clear, and not too repetitive – 10

7. Has a clearly identifiable conclusion – 10

8. Has a convincing argument that draws together the evidence presented – 13

Total: 98 (A)

Style and English Usage – The paper

1. Is clear, direct, and appropriate in tone – 20

2. Is free of misspellings – 10

3. Is free of grammatical mistakes – 10

4. Avoids incomplete sentences – 10

5. Avoids run-on or overly complicated sentences – 10

6. Uses correct punctuation – 10

7. Is free of jargon, cliches, and misused words – 10

8. Cites quotations and/or references correctly – 10

Total: 100 (A)

Comments: Dan, a good improvement over the first version!